Noah's ark was 300 cubits long, 50 cubits wide, and 30 cubits high. Assuming that a cubit is 0.5m, express the dimensions of Noah's ark in m and find the volume in m^3
Answer #1

Length L = 300 cubit = 150 m

Breadth B = 50 cubit = 25 m

Height H = 30 cubit = 15 m

Volume of the Noah's Ark


\small V = L\times B\times H = 150 m\times 25 m \times 15 m = 56250 m^{3}